OpenStudy (anonymous):

A barrel contains 159 gallons of paint and is being drained at a constant rate of 7 gallons per day. Write an equation that models the number of gallons, g after t: days.

OpenStudy (anonymous):

Think of the equation y= mx + b where "b" is your y-intercept, the # of gallons to start with, "x" as the # of days, and "m" as the # of gallons drained off each day. Would you be able to construct an equation by filling in values?

OpenStudy (anonymous):

y=159x+7 ?

OpenStudy (anonymous):

"b" would be the y-intercept, # of gallons to start with and is at the "end" of the equation. "m" is the "drain rate" of 7, and here it is negative because you are decreasing each day.

OpenStudy (anonymous):

If this helps, each day you have 7 less gallons than the previous day, so on the first day, you have "b", then on the second day, you have -7 + b, then -7(2) + b, -7(3) + b, and so on. "b" would be 159, the # you start with for your # of gallons.

OpenStudy (anonymous):

y=-7+159

OpenStudy (anonymous):

g=159-7t

OpenStudy (anonymous):

Just put a "t" after your -7 and you're good to go.